OpenStudy (anonymous):

what are the solutions to 6x^2 = 18x

OpenStudy (mathstudent55):

Subtract 18x to left side. Then divide both sides by 6. Then factor out x. Then set both factors equal to zero and solve each one.

OpenStudy (anonymous):

How am I suppose to divide 6 and 0?

OpenStudy (zehanz):

This is an equation of the form \(a\cdot b=a\cdot c\) . Such an equation can be split up in two parts: 1. \(a=0\) (if a=0, then you get the statement 0=0, which is true) 2. \(b=c\) (if a is not zero, divide both sides by a, leaving b=c). Here you can split up in \(6x \cdot x=6x \cdot 3\), so a=6x, b=x and c=3. So the equation becomes: 1. \(6x=0\) Not difficult to solve... 2. \(x=3\) Hey, a solution!

OpenStudy (mathstudent55):

Follow me. Doone step at a time and show your work. 6x^2 = 18x Subtract 18x from both sides.

OpenStudy (anonymous):

6x^2 - 18 = 0

OpenStudy (mathstudent55):

Don't lose the x with the 18x

OpenStudy (mathstudent55):

It shoudl readd 6x^2 - 18x = 0 Now divide both sides by 6. Whatdo you get?